. Yamoah has an income (M) ofGH¢200 to spend on two commodities Maize and Yam. 

Maize costs GH¢ 1.20 per unit and Yam costs GH¢ 1.1 0 per unit. 


(a) 

Write down Mrs. Yamoah's budget equation. 

(b) 

Write down Mrs. Yamoah's budget line. 

(c) 

What is Mrs. Yamoah's Marginal Rate of Substitution of Maize to Yam? 



1
Expert's answer
2021-06-02T12:26:59-0400

Solution:

a.). Yamoah’s budget equation:

M = PxX + PyY

Where: M = Income = 200

            Px = Price of Maize = 1.20

            Py = Price of Yam = 1.10

b.). Budget line: 200 = 1.20X + 1.10Y

 

c.). MRSx,y = "\frac{MU_{x}}{MU_{y}}"


"\frac{MU_{x}}{P_{x}} = \frac{MU_{y}}{P_{y}}"

 

MUx = "\frac{\partial U} {\partial x} = \frac{1.20} {x}"

 

MUy = "\frac{\partial U} {\partial y} = \frac{1.10} {y}"

 

MRSx,y = "\frac{\partial U} {\partial x} = \frac{1.20} {x}"


MRSx,y = "\frac{\partial U} {\partial y} = \frac{1.10} {x}"


MRSx,y = "\frac{1.09y} { x}"